# Stochastic Dominance Constrained Optimization with S-shaped Utilities: Poor-Performance-Region Algorithm and Neural Network

### v1

@@ Abstract
We investigate the static portfolio selection problem of S-shaped and non-concave utility maximization under first-order and second-order stochastic dominance (SD) constraints. In many S-shaped utility optimization problems, one should require a liquidation boundary to guarantee the existence of a finite concave envelope function. A first-order SD (FSD) constraint can replace this requirement and provide an alternative for risk management. We explicitly solve the optimal solution under a general S-shaped utility function with a first-order stochastic dominance constraint. However, the second-order SD (SSD) constrained problem under non-concave utilities is difficult to solve analytically due to the invalidity of Sion’s maxmin theorem. For this sake, we propose a numerical algorithm to obtain a plausible and sub-optimal solution for general non-concave utilities. The key idea is to detect the poor performance region with respect to the SSD constraints, characterize its structure and modify the distribution on that region to obtain (sub-)optimality. A key financial insight is that the decision maker should follow the SD constraint on the poor performance scenario while conducting the unconstrained optimal strategy otherwise. We provide numerical experiments to show that our algorithm effectively finds a sub-optimal solution in many cases. Finally, we develop an algorithm-guided piecewise-neural-network framework to learn the solution of the SSD problem, which demonstrates accelerated convergence compared to standard neural network approaches.
Keywords: Non-concave utility, portfolio selection, risk constraints, first-order stochastic dominance (FSD), second-order stochastic dominance (SSD), numerical method, neural network
@@ 1 Introduction
S-shaped utility functions, formalized in cumulative prospect theory (Tversky and Kahneman (1992)) and surveyed in behavioral finance (Barberis and Thaler (2003)), capture two salient features of investor behavior: risk seeking in losses and risk aversion in gains. In portfolio selection problems, this non-concavity can induce aggressive tail-risk taking (Carpenter (2000); He and Kou (2018); Liang and Liu (2020)). In particular, when the pricing kernel has a heavy right tail, unconstrained S-shaped optimization typically generates a heavy left tail in terminal wealth, leading to large probabilities of extreme losses, which are the risks that classical VaR/ES-type constraints do not reliably mitigate.
We adopt stochastic dominance (SD) constraints as an implementable approach to control such downside risk under non-concave utilities. First-order SD (FSD) relative to a benchmark wealth $X_{0}$ enforces a quantile-by-quantile floor, offering a distribution-level safety guarantee that often aligns better with practice than hard liquidation boundaries. Second-order SD (SSD) controls cumulative quantiles, providing a flexible, model-free form of downside protection from the pathological risk-seeking induced by the convex (loss) region of S-shaped preferences.
Analytically, SD-constrained utility maximization is well understood for strictly concave utilities via quantile reformulations, duality, and saddle-point arguments (Föllmer and Schied (2016); Wang and Xia (2021)). However, extending these tools to non-concave utilities under SSD is challenging. Sion’s max-min theorem no longer applies in the key saddle-point step, concavification is not guaranteed to be valid because optimizers can fall in regions where the utility and its concave envelope differ (Liang and Liu (2024)), and closed-form solutions are scarce beyond special, technically constrained settings (Wang and Xia (2021)). This paper contributes on three aspects.
@@
Third, we develop an algorithm-guided piecewise-neural-network framework that embeds the PPRA-derived partition and analytic priors into the architecture. This design drastically narrows the functional search space, accelerates convergence, satisfies budget and SSD constraints more quickly, and attains higher objective values than a standard monolithic network, especially in non-concave settings where regular training struggles with infeasibility and local minima.
Methodologically, our approach combines duality and concavification insights (Karatzas et al. (1987); Carpenter (2000); Liang and Liu (2020)) with the quantile formulation of utility maximization (He and Zhou (2011); Föllmer and Schied (2016)) and SD theory (Wang and Xia (2021); Wang, Wei and Xia (2024)). For FSD, we provide a closed-form solution without a liquidation boundary and interpret FSD as a distributional safety floor. For SSD, we translate feasibility into integral inequalities in the quantile domain and design a correction that “repairs” exactly where the unconstrained policy underperforms the benchmark.
@@
Financial implications are immediate. First, FSD guarantees that all terminal-wealth quantiles exceed those of a benchmark, giving a realistic and implementable floor. Second, the optimal/near-optimal policy is to adhere to the benchmark in bad states and follow the classical policy otherwise, yielding transparent risk management. Third, SD constraints significantly reduce the probability of extreme losses induced by S-shaped preferences, beyond what standard VaR/ES controls typically achieve.
@@ 2 Model Settings
where $U:\mathbb{R}\to\mathbb{R}$ is a utility function to be specified in the following different sections. The constraint is called the budget constraint. For a strictly concave utility, the solution of Problem (1) is given by
@@
where the conjugate function $I:(0,\infty)\to\mathbb{R}$ is given by $I(y)\triangleq\arg\sup_{x\in\mathbb{R}}\left[U(x)-xy\right],~y>0$ (We will revisit the definition if $U$ is non-concave) and $\lambda_{\text{cla}}>0$ is a Lagrange multiplier solved from $\mathbb{E}[\rho I(\lambda_{\text{cla}}\rho)]=\overline{x}.$ The problem (1) can be seen as the terminal wealth optimization of the classic continuous-time Merton’s problem in a complete market; see Appendix A of Liang and Liu (2024) for details. In the classic Merton’s problem, the utility function is chosen as a smooth and strictly concave function, including power/log (CRRA, constant relative risk aversion) or exponential (CARA, constant absolute risk aversion) functions.
@@ 3 FSD Problem and Analytical Solution
    $U$ is increasing on $\mathbb{R}$, $U=U_{1}$ is strictly concave on $(B,\infty)$, and $U=U_{2}$ is convex on $(-\infty,B)$.
@@
The conjugate function $I:(0,\infty)\rightarrow\mathbb{R}$ is given by $I(y)=\arg\sup_{x\geqslant B}\left[U(x)-xy\right]=(U_{1}^{\prime})^{-1}(y)$. Note that in Definition 2, there is no requirement for a finite left endpoint of the domain of the utility function (known as the liquidation boundary). We solve the FSD problem (3) with a general S-shaped utility and hence illustrate that using the FSD constraint can replace the liquidation boundary for risk management.
@@
As the objective in Problem (3) is an increasing function of $X$, the optimal wealth $X_{\text{FSD}}^{*}$ of Problem (3) is counter-comonotonic to the pricing kernel $\rho$ (see, e.g., He and Zhou (2011)). Denote by $\xi$ the uniform transformation of $\rho$ such that $\xi$ has the uniform distribution on $[0,1]$ and $Q_{\rho}(\xi)=\rho$. Hence, the optimal wealth $X_{\text{FSD}}^{*}$ under the first-order stochastic dominance constraint should satisfy
$X_{\text{FSD}}^{*}\geqslant Q_{0}(1-\xi),~~Q\in\mathcal{Q}.$
@@
where $\lambda>0$ is a to-be-determined Lagrange multiplier such that $\mathbb{E}[\rho X]=\overline{x}$. To solve Problem (9), we have the following two cases:
@@
In the literature, a liquidation boundary is needed for the optimization of the S-shaped utility, otherwise the problem has no solution (mathematically, it is because one cannot establish a concave envelope for the S-shaped utility without a lower bound in the domain). In the problem of FSD, we do not require the liquidation boundary for the S-shaped utility. The solution is twofold. In some good scenarios, it behaves like the classic solution. In some bad scenarios, it behaves like the benchmark quantile. From the solution, we can see that the FSD constraint acts as a good substitute of the liquidation boundary.
@@ 4.1 SSD Problem under Non-concavity: Analytical Difficulty
Theorem 2 (Theorem 5.10 of Wang and Xia (2021)).
@@
where $y_{\text{SSD}}^{*}:[0,1]\to[0,\infty)$ is a function given by the system
@@
In view of quantile formulation approach, we are going to change the optimization among the random variable $X$ to its quantile function $Q_{X}$. We can hence express the objective in SSD Problem (4) as $\mathbb{E}[U(X)]=\int_{0}^{1}U(Q_{X}(s))\mathrm{d}s.$ Based on the counter-monotonic dependence between the optimal solution $X$ and the pricing kernel $\rho$ (see He and Zhou (2011) and the proof of Theorem 1), we can express the budget constraint as $\mathbb{E}[\rho X]=\int_{0}^{1}Q_{X}(s)Q_{\rho}(1-s)\mathrm{d}s.$ According to Föllmer and Schied (2016), SSD Problem (4) reads as an optimal quantile problem
@@
$\max_{Q\in\mathcal{Q}}\int_{0}^{1}U(Q(s))\mathrm{d}s~~\text{ s.t. }\left\{\begin{aligned} &\int_{0}^{1}Q(s)Q_{\rho}(1-s)\mathrm{d}s\leqslant x,\\ &\inf_{w\in\mathcal{W}_{1}}\left(\int_{[0,1]}(Q(s)-Q_{0}(s))\mathrm{d}w(s)\right)\geqslant 0.\end{aligned}\right$. (15)
@@
Unfortunately, the solution procedure of the SSD Problem under non-concave utility is stuck at this step, because the desired Sion’s max-min theorem requires that $K$ is concave in $Q$, which does not hold generally. In an alternative clue of concavifying $U$ (see Liang and Liu (2020)), one cannot guarantee that the concavification principle is valid (i.e., the optimal wealth variable under the concave envelope is almost surely not located in the region where the original utility and its concave envelope do not coincide). Even if a similar form of Theorem 2 is established for general non-concave utilities, one needs to solve the optimal pair $(y^{*},z^{*})$ from the system (11), which is an infinite-dimensional optimization problem over the functional space. Wang and Xia (2021) propose some explicit optimal solutions based on specific and technical assumptions on $U$, $Q_{\rho}$ and $Q_{0}$. Beyond these, there is no general analytical expression for optimal solutions.
@@
Here for the non-concave utility $U$, we define $I(y)\triangleq\inf\{\arg\sup_{x\in\mathbb{R}}\left[U(x)-xy\right]\},~y>0$ whenever applicable. For example, if $U$ is a S-shaped utility with the domain $L,\infty)$ where $L\in\mathbb{R}$, we have $I(\cdot)$ is well-defined, while if the domain is $\mathbb{R}$, we have $I(\cdot)\equiv$-$\infty$.
Remark: ”Whenever applicable” means that the function $I$ is finite for any $y\in(0,\infty)$. That is, the utility $U$ has a finite concave envelope function (the smallest concave function dominating $U$; see Liang and Liu (2020)). In this case, $I(\cdot)$ is right-continuous and decreasing on $(0,+\infty)$.
In the set $C$, we compare the unconstrained classic solution $I(\lambda Q_{\rho}(t))$ with the SSD benchmark $Q_{0}(1-t)$. If $C=\emptyset$, then it means that the unconstrained solution automatically satisfies the SSD constraint and $Q_{\text{SSD}}^{*}$ should be the same as the unconstrained solution. Next, we discuss the non-trivial case (i.e., $C\neq\emptyset$). For the scenario $t$ on this region, the unconstrained solution $I(\lambda Q_{\rho}(t))$ is smaller (i.e., performing worse) than the SSD benchmark $Q_{0}(1-t)$. From Definition 1, the SSD constraint $Q_{\text{SSD}}^{*}\succeq_{(2)}Q_{0}$ reads as
@@
$z_{\text{SSD}}^{*}(s)=-\int_{s}^{1}\left(Q_{\text{SSD}}^{*}(1-t)-Q_{0}(1-t)\right)\mathrm{d}t\leqslant 0~~\text{ for any }s\in(0,1).$
@@
$z_{\text{SSD}}^{*}(s)=-\int_{s}^{1}\left(I(\lambda(Q_{\rho}(t)-y_{\text{SSD}}^{*}(t)))-Q_{0}(1-t)\right)\mathrm{d}t\leqslant 0~~\text{ for any }s\in(0,1).$ (17)
@@ 4.2 SSD Problem: Numerical Algorithm
1:We solve the classic problem without the SSD constraint and obtain the optimal quantile $Q_{\text{cla}}$. The Lagrange multiplier is denoted by $\lambda_{\text{cla}}\in(0,\infty)$, which is solved from the following equation
@@
$\int_{t}^{1}\left(I(\lambda_{\text{cla}}Q_{\rho}(s))-Q_{0}(1-s)\right)\mathrm{d}s\leqslant 0$
holds for any $t\in[0,1]$, then the optimal solution is $Q_{\text{cla}}(\cdot)\triangleq I(\lambda_{\text{cla}}Q_{\rho}(1-\cdot))$. Otherwise, we start the procedure below.
2:The Lagrange multiplier is initially set as the above $\lambda$ (to be determined at last). Compute the set
@@
$g_{i}(s)=-\int_{s}^{a_{i+1}}\left(I\left(\lambda\left(Q_{\rho}(t)-y_{0}(s)\right)\right)-Q_{0}(1-t)\right)\mathrm{d}t,~~s\in(a_{i},b_{i}).$ (19)
@@
$z_{\text{sub}}(s)\triangleq-\int_{s}^{1}\left(I(\lambda(Q_{\rho}(t)-y_{\text{sub}}(t)))-Q_{0}(1-t)\right)\mathrm{d}t,~s\in[a_{i},a_{i+1}].$ (22)
7:Set $t_{0}=0$ and $y_{0}(t_{0})=0$. After the iteration, we have Eq. (21) for $i=n,\dots,1$ and $y_{\text{sub}}(\cdot)\equiv y_{0}(t_{0})$ on $(t_{0},a_{1}]$. We then verify whether $y_{\textbf{sub}}(\cdot)$ satisfies the monotonicity condition (non-decreasing over $(0,1)$). If the condition holds, proceed to Step 14; otherwise, apply the correction procedure and proceed to Step 8.
8:For $i=n,(n-1),\dots,2$ (Steps 8-12), check whether $y_{0}(\cdot)$ is increasing (non-decreasing) over $(a_{i-1},a_{i+1})$. If yes, skip and proceed to next iteration; if not, proceed to Step 11.
@@
$g_{i}(s)=-\int_{s}^{\bar{s}}\left(I\left(\lambda\left(Q_{\rho}(t)-y_{0}(s)\right)\right)-Q_{0}(1-t)\right)\mathrm{d}t,~~s\in(a_{i-1},b_{i-1}),$
where $\bar{s}=\inf\{t\geqslant a_{i}|y_{0}(t)-y_{0}(s)\geqslant 0\},t\in[a_{i},b_{i}).$
@@
$z^{*}_{i}(t)=-\int_{t}^{1}\left(I\left(\lambda\left(Q_{\rho}(s)-y_{0}(t_{i})\right)\right)-Q_{0}(1-s)\right)\mathrm{d}s,~~t\in(\max\{a_{i},t_{i}\},a_{i+1}).$
If for any $i=1,\dots,n$, $z^{*}_{i}(\cdot)\leqslant 0$ always holds on the interval $(\max\{a_{i},t_{i}\},a_{i+1})$, this $Q_{\text{sub}}$ is sub-optimal.
@@
For some very poorly-performed scenarios $t\in C$, the function $y_{0}$ is adopted such that $Q_{\text{sub}}(1-t)=I(\lambda(Q_{\rho}(t)-y_{0}(t)))=Q_{0}(1-t)$.
The general idea in the construction of the sub-optimal solution to SSD Problem (4) is: When $z_{\text{sub}}\leqslant 0$, we use the classic solution to achieve optimality (now $y_{\text{sub}}=0$ or constant, and $Q_{\text{SSD}}$ is the form of $I$); When $z_{\text{sub}}>0$, we set $y_{\text{sub}}=y_{0}$ such that $Q_{\text{sub}}=Q_{0}$ to satisfy the SSD constraint. Throughout the algorithm design, we need to guarantee that $y_{\text{sub}}$ is increasing and non-negative.
We therefore propose Algorithm 1 below, named the Poor-Performance-Region Algorithm. Based on Theorem 2, we design a closed-form sub-optimal solution to Problem (14):
@@
    In Step 2: In the whole procedure, we are solving out the structure of the optimal solution and the Lagrange multiplier. For the latter, note that the initial Lagrange multiplier may not satisfy the budget constraint. But it is a good initial point to start the algorithm. It will be determined in the last step.
@@
    In Steps 4-5: For any $t\in[b_{i},a_{i+1})$, we have I(λ^* (Q_ρ(s) - y_0(t) ) ) - Q_0(1-s) ⩾I(λ^* (Q_ρ(s) ) ) - Q_0(1-s) ⩾0,    s ∈(t, a_i+1]. We then compute g_i(t) = - ∫_t^a_i+1 ( I(λ^* (Q_ρ(s) - y_0(t) ) ) - Q_0(1-s) ) ds ⩽0. Hence, any $t\in[b_{i},a_{i+1})$ satisfies the SSD constraint. We desire to search the first point $t_{i}$ on $(a_{i},b_{1})$ which does not satisfy the constraint.

### v2

@@ Abstract
We investigate the static portfolio selection problem of S-shaped (non-concave) utility maximization under first-order and second-order stochastic dominance (SD) constraints. In many S-shaped utility optimization problems, one should require a liquidation boundary to guarantee the existence of a finite concave envelope function. A first-order SD (FSD) constraint can replace this requirement and provide an alternative for risk management. We explicitly solve the optimal solution under a general S-shaped utility function with a first-order stochastic dominance constraint. However, the second-order SD (SSD) constrained problem under non-concave utilities is difficult to solve analytically due to the invalidity of Sion’s max-min theorem. For this reason, we propose a numerical algorithm to obtain a plausible and sub-optimal solution for general non-concave utilities. The key idea is to detect the poor performance region with respect to the SSD constraints, characterize its structure and modify the distribution on that region to obtain (sub-)optimality. A key financial insight is that the decision maker should follow the SD constraint on the poor performance scenario while conducting the unconstrained optimal strategy otherwise. We provide numerical experiments to show that our algorithm effectively finds a sub-optimal solution in many cases. Finally, we develop an algorithm-guided piecewise-neural-network framework to learn the solution of the SSD problem, which demonstrates accelerated convergence compared to standard neural network approaches.
Keywords: non-concave utility, portfolio selection, risk constraints, first-order stochastic dominance (FSD), second-order stochastic dominance (SSD), poor performance region, numerical algorithm, neural network
@@ 1 Introduction
S-shaped utility functions, formalized in cumulative prospect theory (Tversky and Kahneman (1992)) and surveyed in behavioral finance (Barberis and Thaler (2003)), capture two salient features of the investor behavior: risk seeking in losses and risk aversion in gains. In portfolio selection problems, this non-concavity can induce aggressive tail-risk taking (Carpenter (2000); He and Kou (2018); Chen, Hieber and Nguyen (2019); Dong and Zheng (2020); Dai et al. (2022); Liang, Liu and Zhang (2025)). In particular, when the pricing kernel has a heavy right tail, unconstrained S-shaped optimization typically generates a heavy left tail in terminal wealth, leading to large probabilities of extreme losses, which are the risks that classical VaR/ES-type constraints do not reliably mitigate.
In the field of optimization, stochastic dominance (SD) constraints are widely applied to portfolio selection problems with risk management (Rockafellar (1970); Dentcheva and Ruszczyński (2003, 2006); Wang and Xia (2021); Wang, Wei and Xia (2024)). In this paper, we adopt stochastic dominance constraints as an implementable approach to control such downside risk under non-concave utilities. First-order SD (FSD) relative to a benchmark wealth $X_{0}$ enforces a quantile-by-quantile floor, offering a distribution-level safety guarantee that often aligns better with practice than hard liquidation boundaries. Second-order SD (SSD) restricts cumulative quantiles, providing a flexible, model-free form of downside protection from the pathological risk-seeking induced by the convex (loss) region of S-shaped preferences.
Analytically, SD-constrained utility maximization is well investigated for strictly concave utilities via quantile reformulations, duality, and saddle-point arguments (Wang and Xia (2021)). However, extending these tools to non-concave utilities under SSD is challenging. Sion’s max-min theorem no longer applies in the key saddle-point step, and the concavification principle is not guaranteed to be valid because optimizers can fall in regions where the utility and its concave envelope differ (e.g., Bichuch and Sturm (2014); Dai et al. (2022); Pesenti, Wang and Wang (2025)), and closed-form solutions are scarce beyond special and technically constrained settings (Wang and Xia (2021)). This paper contributes in three aspects.
@@
Third, we develop an algorithm-guided piecewise-neural-network framework that embeds the PPRA-derived partition and analytic priors into the architecture. This design drastically narrows the functional search space, accelerates convergence, satisfies budget and SSD constraints more quickly, and attains higher objective values than a standard monolithic network, especially in non-concave settings where regular training struggles with infeasibility and local minima (Scarselli and Tsoi (1998)).
Methodologically, our approach combines duality and concavification insights (Karatzas et al. (1987); Carpenter (2000); Liang and Liu (2020)) with the quantile formulation of utility maximization (He and Zhou (2011); Xu (2016); Ghossoub and Zhu (2025)) and SD constrained optimization (Wang and Xia (2021); Wang, Wei and Xia (2024)). For FSD, we provide a closed-form solution without a liquidation boundary and interpret FSD as a distributional safety floor. For SSD, we translate feasibility into integral inequalities in the quantile domain and design a correction that “repairs” exactly where the unconstrained policy underperforms the benchmark.
@@
Financial implications are immediate. First, FSD guarantees that all terminal-wealth quantiles exceed those of a benchmark, giving a realistic and implementable floor. Second, the optimal/near-optimal policy is to adhere to the benchmark in bad states and follow the classical policy otherwise, yielding transparent risk management. Third, SD constraints significantly reduce the probability of extreme losses induced by S-shaped preferences, beyond what standard VaR/ES constraints typically achieve.
@@ 2 Model Settings
where $U:\mathbb{R}\to\mathbb{R}$ is a utility function to be specified in the following sections. The constraint is called the budget constraint. For a strictly concave utility $U$, the solution of Problem (1) is given by
@@
where the conjugate function $I:(0,\infty)\to\mathbb{R}$ is given by $I(y)\triangleq\arg\sup_{x\in\mathbb{R}}\left[U(x)-xy\right],~y>0$ (we will revisit the definition if $U$ is non strictly concave) and $\lambda_{\text{cla}}>0$ is a Lagrange multiplier solved from $\mathbb{E}[\rho I(\lambda_{\text{cla}}\rho)]=\overline{x}.$ The problem (1) can be seen as the terminal wealth optimization of the classic continuous-time Merton’s problem in a complete market; see Appendix A of Liang and Liu (2024) for details. In the classic Merton’s problem, the utility function is chosen as a smooth and strictly concave function, including power/log (CRRA, constant relative risk aversion) or exponential (CARA, constant absolute risk aversion) functions.
@@ 3 FSD Problem and Analytical Solution
    $U$ is increasing on $\mathbb{R}$, $U=U_{1}$ is differentiable and strictly concave on $(B,\infty)$, and $U=U_{2}$ is convex on $(-\infty,B)$.
@@
Note that in Definition 2, there is no requirement for a finite left endpoint of the domain of the utility function (known as the liquidation boundary in He and Kou (2018)). We solve the FSD problem (3) with a general S-shaped utility and hence illustrate that using the FSD constraint can replace the liquidation boundary for risk management.
@@
As the objective in Problem (3) is an increasing function of $X$, the optimal wealth $X_{\text{FSD}}^{*}$ of Problem (3) is counter-monotonic to the pricing kernel $\rho$ (see, e.g., He and Zhou (2011); Wei (2018)). Denote by $\xi$ the quantile transformation of $\rho$ such that $\xi$ has the uniform distribution on $[0,1]$ and $Q_{\rho}(\xi)=\rho$. Hence, the optimal wealth $X_{\text{FSD}}^{*}$ under the first-order stochastic dominance constraint should satisfy
$X_{\text{FSD}}^{*}\geqslant Q_{0}(1-\xi).$
@@
where $\lambda>0$ is a Lagrange multiplier to be determined such that $\mathbb{E}[\rho X]=\overline{x}$. To solve Problem (9), we have the following two cases:
@@
In the literature, a liquidation boundary is needed for the optimization of the S-shaped utility, otherwise Problem (3) has no solution (mathematically, it is because one cannot establish a concave envelope for the S-shaped utility without a lower bound in the domain). In the FSD problem, we do not require the liquidation boundary for the S-shaped utility. The solution is twofold. In some good scenarios, it behaves like the classic solution. In some bad scenarios, it behaves like the benchmark quantile. From the solution, we can see that the FSD constraint acts as a good substitute for the liquidation boundary.
@@ 4.1 SSD Problem under Non-concavity: Analytical Difficulty
Lemma 1 (Theorem 5.10 of Wang and Xia (2021)).
@@
where $y_{\text{SSD}}^{*}:(0,1)\to[0,\infty)$ is a function given by the system
@@
In view of the quantile formulation approach, we are going to reformulate the optimization over the random variable $X$ in terms of its quantile function $Q_{X}$. We can hence express the objective in SSD Problem (4) as $\mathbb{E}[U(X)]=\int_{0}^{1}U(Q_{X}(s))\mathrm{d}s.$ Based on the counter-monotonic dependence between the optimal solution $X$ and the pricing kernel $\rho$ (see He and Zhou (2011) and the proof of Theorem 1), we can express the budget constraint as $\mathbb{E}[\rho X]=\int_{0}^{1}Q_{X}(s)Q_{\rho}(1-s)\mathrm{d}s.$ According to Föllmer and Schied (2016), SSD Problem (4) reads as an optimal quantile problem
@@
$\max_{Q\in\mathcal{Q}_{2}(Q_{0})}\int_{0}^{1}U(Q(s))\mathrm{d}s~~\text{ s.t. }\left\{\begin{aligned} &\int_{0}^{1}Q(s)Q_{\rho}(1-s)\mathrm{d}s\leqslant\bar{x},\\ &\inf_{w\in\mathcal{W}_{1}}\left(\int_{[0,1]}(Q(s)-Q_{0}(s))\mathrm{d}w(s)\right)\geqslant 0.\end{aligned}\right$. (15)
@@
Unfortunately, the solution procedure of the SSD Problem under the non-concave utility is stuck at this step, because the desired Sion’s max-min theorem requires that $K$ is concave in $Q$, which does not hold generally. In an alternative clue of concavifying $U$ (see Liang and Liu (2020)), one cannot guarantee that the concavification principle is valid (i.e., the optimal wealth variable under the concave envelope is almost surely not located in the region where the original utility and its concave envelope do not coincide); here, the concave envelope is defined as the smallest concave function dominating $U$. Even if a similar form of Lemma 1 is established for general non-concave utilities, one needs to solve the optimal pair $(y^{*},z^{*})$ from the system (11), which is an infinite-dimensional optimization problem over the functional space. Wang and Xia (2021) propose some explicit optimal solutions based on specific and technical assumptions on $U$, $Q_{\rho}$ and $Q_{0}$. Beyond these, there is no general analytical expression for optimal solutions.
@@
Here for the non-concave utility $U$, we define $I(y)\triangleq\inf\{\arg\sup_{x\in\mathbb{R}}\left[U(x)-xy\right]\},~y>0$ whenever applicable.
Remark 1.
“Whenever applicable” means that the function $I(y)$ is finite for any $y\in(0,\infty)$. That is, the utility $U$ has a finite concave envelope function. In this case, $I(\cdot)$ is right-continuous and decreasing on $(0,\infty)$. For example, if $U$ is an S-shaped utility with the domain $L,\infty)$ where $L\in\mathbb{R}$, we have $I(\cdot)$ is well-defined, while if the domain is $\mathbb{R}$, $I(\cdot)\equiv$-$\infty$.
In the set $C$, we compare the unconstrained classic solution $I(\lambda Q_{\rho}(t))$ with the SSD benchmark $Q_{0}(1-t)$. If $C=\emptyset$, then it means that the unconstrained solution automatically satisfies the SSD constraint, and $Q_{\text{SSD}}^{*}$ should be the same as the unconstrained solution. Next, we discuss the non-trivial case (i.e., $C\neq\emptyset$). For scenarios $t$ on this region, the unconstrained solution $I(\lambda Q_{\rho}(t))$ is smaller (i.e., performing worse) than the SSD benchmark $Q_{0}(1-t)$. From Definition 1, the SSD constraint $Q_{\text{SSD}}^{*}\succeq_{(2)}Q_{0}$ reads as
@@
$z_{\text{SSD}}^{*}(s)=-\int_{s}^{1}\left(Q_{\text{SSD}}^{*}(1-t)-Q_{0}(1-t)\right)\mathrm{d}t\leqslant 0~~\text{ for any }s\in(0,1),$
@@
$z_{\text{SSD}}^{*}(s)=-\int_{s}^{1}\left[I(\lambda(Q_{\rho}(t)-y_{\text{SSD}}^{*}(t)))-Q_{0}(1-t)\right]\mathrm{d}t\leqslant 0~~\text{ for any }s\in(0,1).$ (17)
@@ 4.2 SSD Problem: Numerical Algorithm
1:We solve the classic problem without the SSD constraint and obtain the optimal quantile $Q_{\text{cla}}(\cdot)\triangleq I(\lambda_{\text{cla}}Q_{\rho}(1-\cdot))$. The Lagrange multiplier is denoted by $\lambda_{\text{cla}}\in(0,\infty)$, which is solved from the following equation
@@
$\int_{t}^{1}\left[I(\lambda_{\text{cla}}Q_{\rho}(s))-Q_{0}(1-s)\right]\mathrm{d}s\leqslant 0$
holds for any $t\in[0,1]$, then the optimal solution is $Q_{\text{cla}}(\cdot)$. Otherwise, we start the procedure below.
2:The Lagrange multiplier $\lambda$ is initially set as the above (to be determined at last). Compute the set
@@
$g_{i}(s)=-\int_{s}^{a_{i+1}}\left[I\left(\lambda\left(Q_{\rho}(t)-y_{0}(s)\right)\right)-Q_{0}(1-t)\right]\mathrm{d}t,~~s\in(a_{i},b_{i}).$ (19)
@@
$z_{\text{sub}}(s)\triangleq-\int_{s}^{1}\left[I(\lambda(Q_{\rho}(t)-y_{\text{sub}}(t)))-Q_{0}(1-t)\right]\mathrm{d}t,~s\in[a_{i},a_{i+1}].$ (22)
7:Set $t_{0}=0$ and $y_{0}(t_{0})=0$. After the iteration, we have Eq. (21) for $i=n,\dots,1$ and $y_{\text{sub}}(\cdot)\equiv y_{0}(t_{0})$ on $(t_{0},a_{1}]$. We then verify whether $y_{\text{sub}}(\cdot)$ satisfies the monotonicity condition (non-decreasing over $(0,1)$). If the condition holds, proceed to Step 14; otherwise, apply the correction procedure and proceed to Step 8.
8:For $i=(n-1),\dots,2$ (Steps 8-12), check whether $y_{0}(\cdot)$ is increasing (non-decreasing) over $(a_{i-1},a_{i+1})$. If yes, skip and proceed to next iteration; if not, proceed to Step 9.
@@
$g_{i}(s)=-\int_{s}^{\bar{s}}\left[I\left(\lambda\left(Q_{\rho}(t)-y_{0}(s)\right)\right)-Q_{0}(1-t)\right]\mathrm{d}t,~~s\in(a_{i-1},b_{i-1}),$
where $\bar{s}=\inf\{u\in[a_{i},b_{i})|y_{0}(u)-y_{0}(s)\geqslant 0\}.$
@@
$z^{*}_{i}(t)=-\int_{t}^{1}\left[I\left(\lambda\left(Q_{\rho}(s)-y_{0}(t_{i})\right)\right)-Q_{0}(1-s)\right]\mathrm{d}s,~~t\in(\max\{a_{i},t_{i}\},a_{i+1}).$
If for any $i=1,\dots,n$, $z^{*}_{i}(\cdot)\leqslant 0$ always holds on the interval $(\max\{a_{i},t_{i}\},a_{i+1})$, this $Q_{\text{sub}}$ is feasible and sub-optimal.
@@
For some very poorly performing scenarios $t\in C$, the function $y_{0}$ is adopted such that $Q_{\text{sub}}(1-t)=I(\lambda(Q_{\rho}(t)-y_{0}(t)))=Q_{0}(1-t)$.
The general idea in the construction of the sub-optimal solution to SSD Problem (4) is: when $z_{\text{sub}}\leqslant 0$, we use the classic solution to achieve optimality (now $y_{\text{sub}}=0$ or constant, and $Q_{\text{SSD}}$ is the form of $I$); when $z_{\text{sub}}>0$, we set $y_{\text{sub}}=y_{0}$ such that $Q_{\text{sub}}=Q_{0}$ to satisfy the SSD constraint. Throughout the algorithm design, we need to guarantee that $y_{\text{sub}}$ is increasing and non-negative. As the optimal structure in Theorem 5.10 of Wang and Xia (2021) comes from a strictly concave utility setting, the sub-optimal solution under our setting of a non-concave utility $U$ is actually the “solution” of the concave envelope of $U$.
We therefore propose Algorithm 1 below, named the Poor-Performance-Region Algorithm (PPRA). Based on Theorem 1, we design a closed-form sub-optimal solution to Problem (14):
@@
    In Step 2: In the whole procedure, we are solving out the structure of the optimal solution and the Lagrange multiplier. For the latter, note that the initial Lagrange multiplier may not satisfy the budget constraint. But it is a good initial point to start the algorithm. It will be determined in the final step.
@@
    In Steps 4-5: For any $t\in[b_{i},a_{i+1})$, we have I(λ^* (Q_ρ(s) - y_0(t) ) ) - Q_0(1-s) ⩾I(λ^* Q_ρ(s) ) - Q_0(1-s) ⩾0,    s ∈(t, a_i+1]. We then compute g_i(t) = - ∫_t^a_i+1 [ I(λ^* (Q_ρ(s) - y_0(t) ) ) - Q_0(1-s) ] ds ⩽0. Hence, any $t\in[b_{i},a_{i+1})$ satisfies the SSD constraint. We desire to search the first point $t_{i}$ on $(a_{i},b_{i})$ which does not satisfy the constraint.